OpenStudy (anonymous):

given sec(theta)=5 find cos(90-theta)

OpenStudy (raden):

first, sec(θ) = 1/sin(θ) so, sec(θ) = 5 1/cos(θ) = 5 cos(θ) = 1/5 then find sin(θ) by using this identity : sin^2 (θ) + cos^2 (θ) = 1 so, sin^2 (θ) + (1/5)^2 = 1 sin^2 (θ) = 1 - 1/25 sin^2 (θ) = 24/25 sin(θ) = sqrt(24/25) =1/5 sqrt(24) = 1/5 * 2sqrt(6) = 2/5 sqrt(6) then use the identity of cos(90-θ) = sin(θ) therefore, cos(90-θ) = sin(θ) = 2/5 sqrt(6)